Identify the intervals in which the function f on R defined by
f(x) = 4x^3-6x^2-72x+15 is both increasing and decreasing ?
1
Expert's answer
2019-10-01T11:18:55-0400

To determine intervals where the function strictly increases or decreases, we need to first find the derivative of the function

"f(x) = 4 x^3 - 6 x^2 - 72 x + 15"


"\frac{d}{dx}f(x)=\frac{d}{dx}(4 x^3 - 6 x^2 - 72 x + 15) = 12 x^2 -12x -72=12( x^2 -x -6)=12(x-3)(x+2)" .


Putting "f'(x)=0."

"12(x-3)(x+2)=0 \\
(x-3)(x+2)=0"

So "x=3" and "x=-2" .

The points divide the real line into 3 disjoint intervals, i.e.

"(-\infty; -2)\bigcup (-2;3)\bigcup (3;\infty)"

"\def\arraystretch{1.5}
 \begin{array}{c|c|c|c}
 \text{ interval} & (-\infty;-2) & (-2;3) &(3;\infty) \\ \hline
 \text{value } x& x<-2 & -2<x<3 & x>3 \\ \hline
\text{sign of } f'(x)=12(x-3)(x+2)& (-)(-)=+>0&(-)(+)=-<0&(+)(+)=+>0\\ \hline
 \text{nature of function } f & \text{ increasing } &\text{ decreasing} & \text{ increasing }
\end{array}"

f(x) is increasing in "(-\infty; -2)\bigcup (3;\infty)"

f(x) is decreasing in "(-2;3)"